Assign costs -- In August , Jay Corp determined its cost per equivalent units in department M were $2.40 per equivalent unit of raw materials and $1.80 per equivalent unit of conversion costs. During the month department M completed and transferred out 97000 units to finished goods inventory. At the month end it has 35,000 of partially completed units. These partially completed units were equal to 15,000 equivalent units of direct materials and 12,000 equivalent units of conversion costs. 1. Determine the totals cost of direct materials and conversion costs that should be assigned to a: Units completed and transferred out b: Units in ending work in process inventory 2. What were the total costs to account for in department M?
Answer #1

1. Finished goods

a. direct material cost = 97000*2.4 = 232800

cost of conversion = 97000*1.8 = 174600

b. Work in progress

direct material cost = 15000*2.4 = 36000

cost of conversion = 12000*1.8 = 21600

2

Total cost = Finished goods material cost + Finished goods cost of conversion + Work in progress direct material cost + Work in progress  cost of conversion = 232800 +174600 +36000 +21600 =

465000